OpenStudy (anonymous):

how do you graph f(x) = y +12 ?

OpenStudy (anonymous):

there is something weird going on...you are looking for a function of x...but your function is in terms of y...

OpenStudy (anonymous):

yeah, and I don't know how to graph something like that. I have to graph that and g(x) = y^2 and find the area between them. The area is 343/6, but the online submission tells me that my graph is wrong.

OpenStudy (anonymous):

If the function was given as x = y + 12 you could find out that y = x-12 g would be sqrt x = y...

OpenStudy (anonymous):

but is isn't....or can I do it that way, even though it is given in terms of f(x) and g(x)?

OpenStudy (anonymous):

the notation they gave you is bogus so not sure what to try. Really the variables that are used are immaterial - but they should match. What you have been given is jibberish. It basically says given some value for x....y+12 = something. It doesn't matter what x I give you...y + 12 would equal something else.

OpenStudy (anonymous):

okay. I was originally treating the two functions as if they had the variable x in them instead, and the computer gave me an error message. i don't know what to think. Thanks for the help